MOBY Photo Album - 20190717_hawaii-2019-03

20190717_hawaii-2019-03

Hi Steph,

Riley & I managed to collect some BS04cfg101 & RS03cfg001 data yesterday, which are at /ftp1/Mike/HI-2019-03/ = ready for your data processing help.

I wanted to setup both specs to collect data somewhat-simultaneously - I'm not sure that doing one-spec-at-a-time was such a bad idea now...

There are photos of the setup, in the cal van next to the tent. Both specs' bare fiber-bundle fibers were looking into the Spectralon sphere. There were no internal aux data (degC, %RH) collected from the specs.

Day01 had as input to the sphere a Thorlabs LED. I think this was the first use of this source. Carol found this in 2015 and we bought one = LED driver / power supply, aka Thorlabs DC2100, plus 5x different LED "heads". Last night we scanned the M530L3 head, which should be centered ~ 530 nm. The LED pwr sup was set at 9 mA ( == LOW power ), as noted in the log sheet.

I ran the Labsphere Si PhotoDiode monitor at the sphere monitor port #A. This would be the PD that Carol had at NIST for cals (which you should have data from?) but, the detector was not a UV-enhanced kine, so it has no response in the UV, and we will try to re-purchase another PD with a UV-enhanced detector. The PD's electrical-current to voltage pre-amp was the one Carol had at NIST for the NIST cals, aka Femto DLPCA-200. The Voltage output from the pre-amp was logged via the 34970A DVM using software Vdc6_5.vi, which you have seen output from before. Someday I will update this VI for use with a PD monitor... Last night, the PD preamp went to DVM chan#3 (i.e. setup for high precision Voltage measurement, typically for a shunt during Ed/Lu absolute lamp cals), and DVM chan#5 had the Radio Shack 10 kOhm thermistor for ambient temperature monitor. For now the only indication of what gain setting the PD preamp was set at is noted in the log sheet - last night was @ low gain, 10^4. This will someday be included in the log file's header...

The log file from the DVM is in the HI-2019-03/BSRS/ subdir on the NAS. The 1x day01 log sheet is in the HI-2019-03/doc/ subdir on the NAS.

One problem last night: I did not power ON the TEC for the PD monitor. Tonight I will try harder to turn that bugger on.

We had to re-load the Andor SOLIS software onto 2x HP laptops to control the BS & RS. So I'm hoping that I set all acquisition parameters correctly. The data acq sequence was 2x bac, 4x sig, 2x bac.

Tonight we will attempt to collect some Ekspla laser scans via BS & RS with all bare bundle fibers illuminated.

Good Luck with this! MF


20190716_01=day01

20190716_02=day01

20190716_03=day01

20190716_04=day01

20190716_05=day01

20190716_06=day01

20190716_07=day01

20190716_08=day01

20190716_09=day01-Thorlabs-LED

20190716_10=day01-Thorlabs-LED

20190716_11=day01-Thorlabs-LED

20190716_12=day01-Thorlabs-LED

20190716_13=day01-Thorlabs-LED

20190716_14=day01-Thorlabs-LED

20190716_15=day01-Thorlabs-LED

20190716_16=day01-Thorlabs-LED

20190716_17=day01-Labsphere-PD

20190716_18=day01-Labsphere-PD

Stephanie Flora - Moss Landing Marine Laboratories